Systems Staff Meeting Notes 8/3/98

Systems Staff Meeting Notes
August 3, 1998

Present: Rich Bennett, Brian Burton, Bill Covey, Winston Harris, Bobbie Meng, Suzy Shaw, Leota Whitfield

Announcements
  1. Mellon meeting this morning - the Libraries may get another Mellon grant. Need to do a comparison of techniques to do this type of indexing - a combination of OCR & Scan techniques. Three newspapers will be completely indexed. Will meet again in 2 weeks.
  2. Bill distributed a simplified diagram of our place in the campus CORE. The Libraries are very far ahead in technology - parts of the Health Center are still on thinnet, Criser lives on hand-me-downs, still has token ring technology.
  3. Judy Bowen - needs a list of control units & 3287 printers (about ten items) - 3 of which are to be surveyed.
  4. A committee has formed - Winston Harris & Barbara Oliver - to search for Destry's replacement.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. PK Yonge - pick up scheduled for tomorrow. How much they take depends upon how many helpers they bring, what vehicles. The most the van can hold is 10 units, will depend on the number of trips they can make. Lee has 37 units ready to go, color-coded pick ups ready & paperwork is done. Want to make sure Iona knows how much is here waiting for her.
  2. CISCO routers have not been picked up yet.
  3. 46 Cows have been logged in. Six public classroom machines are to go to the public Circulation area, classroom will get 6 new Cows to replace them. MSL gets 28 new Cows plus 3 with zip drives built in. Four machines (out of the old shipment) will go to Maps - they will need more memory - need to go from 32M to 64M.
  4. Have received the 6 little Cows with Windows '98. These will be the 3270 replacements. Need to develop a method to secure them.
  5. MSL picked up 22 Centurion Guards.
  6. Next ordering spree we will get P2's for the Analysts.
  7. Exchange upgrade to take place before we get the new server - COMPAQ - will be 20-55 Gig. Will replace NT2. 128M - should take care of stability problems.
  8. Mac software can go to PK Yonge. AIR software & Apple Internet Router software will have to be surveyed. SNAPS server software may also be surveyed.
Liaisons
  1. Marcia has taken Stacey Marsh's position. She will no longer be a liaison - Jim Stevens & Hang Soo Hoo will now take over in that department.
  2. Bill is working with Trudi Di Trolio on training.
  3. CD access - Women's Resource & Philosopher's Index have been reporting access problems.
Networking
  1. IP management problems - will call CISCO.
  2. Public printing - no complaints. Need to move the backup server to the vault. The SMATHERS Novell server is part of the routing for Smathers_PS (public printing server). If Smathers goes down, call NERDC, have them set SMATHERS_PS to be permitted for GNS service instead of SMATHERS Novell.
  3. The slowdown we experienced last week seems to be OK - repeater swap-out in LAC seems to have fixed it.
  4. NT2 is back in the Network Neighborhood - the work around shortcut works. The Microsoft Win'95 escalation team is trying to reach Bill about this situation.
  5. Query from Rob Roberts - questioning wiring for digital phones in the North Tower - it is not there.
  6. Lee talked to Earl about MSL wiring. Plan to put a junction box & 2 strings where the kiosk is located. They want to move the kiosk, & have put in an extra string. Lee suggested that they pull a new line & put extra wire in since they don't want jumper or wire on the floor.
  7. MSL's furniture is expected by the middle of this month.
Software Projects Status
  1. Not too many software people were here last week.
  2. First FY invoicing - some problems with ILL data being included.
  3. NT4 timeout this morning - took a powder during a tape cartridge change. It loses it's counter & can't rewrite it's primary key.
DP Coordinator
  1. Coming up on the 2 week interval between semesters.
  2. OPS budget is back this year.
  3. The Tape Library A/C was out last week. The compressor kicked in. If the A/C goes down, need to turn the humidistat up to 100%.
Calendar
  1. Sharon out today.
  2. Martha Hruska will be back on August 17.
  3. Lee out August 17-18.
  4. Bill out August 21 (Harris Corporation meeting).
  5. Lee out September 1.
  6. Lee out Sept. 17 - on vacation, then on to retirement.


Back to the Systems Homepage

Last updated February 24, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu